OpenStudy (anonymous):

if a shoe is 24 centimeters how many shoes would it take to make one meter than one lilometer

OpenStudy (anonymous):

What's a lilometer?

OpenStudy (anonymous):

i mean klilometer

OpenStudy (anonymous):

i wrote it byaccident

OpenStudy (anonymous):

Ok, so a shoe is 24 centimeters. How many centimeters in a meter?

OpenStudy (anonymous):

100

OpenStudy (anonymous):

And how many meters in a kilometer?

OpenStudy (anonymous):

1000

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

does it take that long to reply

OpenStudy (anonymous):

Ok so: \[1\ shoe = 24 cm \implies 1 = \frac{1\ shoe}{24cm}\] \[100 cm = 1m \implies 1 = \frac{100cm}{1m}\] \[1000m = 1km \implies 1 = \frac{1000m}{1km}\] \[\implies 1m \cdot \frac{100cm}{1m} \cdot \frac{1\ shoe}{24cm} = \frac{100}{24}shoe\]

OpenStudy (anonymous):

Then you can multiply that by another thousand to find how many you need for a kilometer